find the locus of centroid of tirangle if two vertex are given and third vertex lies on the line

Grade:12th pass

1 Answers

Gopal Gopakumar
7 years ago
Let the given points be (A1,B1) , and (A2,B2) . Write the given line in the form of y = mx + c .
Hence, the third point becomes (x, mx + c) .
Putting in equation of centroid gives,
 
(A1 + A2 + x)/3 = h ( x – coordinate of locus)  
and  (B1 + B2 + mx + c)/3 = k ( y – cordinate of locus)
 
Rearranging the two equation gives you,
 
3h – A1 – A2 = x   and  (3k – B1 – B2 – c)/m = x
 
equating the two and putting h = x and k = y finally will give the required answer.